Activity-Based Supplier Costing

Clearsound uses Alpha Electronics and La Paz Company to buy two electronic components used in the manufacture of its cell phones: Component 125X and Component 30Y. Consider two activities: testing components and reordering components. After the two components are inserted, testing is done to ensure that the two components in the phones are working properly. Reordering occurs because one or both of the components have failed the test and it is necessary to replenish component inventories. Activity cost information and other data needed for supplier costing are as follows:

I. Activity Costs Caused by Suppliers (testing failures and reordering as a result)

Activity Costs    
Testing components $1,200,000
Reordering components 300,000

II. Supplier Data

Alpha Electronics La Paz Company
125X 30Y 125X 30Y
Unit purchase price $10 $26 $12 $28
Units purchased 120,000 68,400 15,000 15,000
Failed tests 1,300 780 10 10
Number of reorders 60 40 0 0

Required:

Determine the cost of each supplier by using ABC. Round Test and Reorder rates to the nearest dollar, and final answers to the nearest cent.

Alpha Electronics La Paz Company
125X 30Y 125X 30Y
Unit cost: $ $ $ $
Answer #1

Activity-based concept states that costs should be recognized on the basis of activities which they can be traced back to and not specific blanket rate.

First of all , cost per each activity needs to be computed

Alpha Electronics

La paz company

Total

Cost per

125X

30Y

125X

30Y

Activity

Cost

activity

Failed tests

1300

780

10

10

2100

1,200,000

571

Number of reorders

60

40

0

0

100

300,000

3000

Each activity rate needs to be multiplied with the number of activities to get total cost in each product as and added to purchase cost to arrive at the cost per unit as:

Alpha Electronics

La Paz company

125X

30Y

125X

30Y

Total

Purchase price

10

26

12

28

Units purchased

120,000

68400

15000

15000

Total purchase price

$    1,200,000

$ 1,778,400

$      180,000

$             420,000

$           3,578,400

Testing components

$        742,857

$     445,714

$           5,714

$                 5,714

$           1,200,000

Reordering components

$        180,000

$     120,000

$               300,000

Total costs (a)

$    2,122,857

$ 2,344,114

$      185,714

$             425,714

$           5,078,400

Units purchased (b)

120,000

68400

15000

15000

Unit cost (a/b)

$             17.69

$          34.27

$           12.38

$                 28.38

742857 = 1300 *571 & 180000 = 60 * 3000 and so on